# Presence or Absence: Are Unknown Word Usages in Dictionaries?

### v1

@@ Abstract
In this work, we outline the components and results of our system submitted to the AXOLOTL-24 shared task for Finnish, Russian and German languages. Our system is fully unsupervised. It leverages a graph-based clustering approach to predict mappings between unknown word usages and dictionary entries for Subtask 1, and generates dictionary-like definitions for those novel word usages through the state-of-the-art Large Language Models such as GPT-4 and LLaMA-3 for Subtask 2. In Subtask 1, our system outperforms the baseline system by a large margin, and it offers interpretability for the mapping results by distinguishing between matched and unmatched (novel) word usages through our graph-based clustering approach. Our system ranks first in Finnish and German, and ranks second in Russian on the Subtask 2 test-phase leaderboard. These results show the potential of our system in managing dictionary entries, particularly for updating dictionaries to include novel sense entries. Our code and data are made publicly available [^1] https://github.com/xiaohemaikoo/axolotl24-ABDN-NLP.
@@ 1 Introduction
  • For Subtask 1, we keep the workflow of the AXOLOTL-24 baseline system unchanged, which includes three components: producing embeddings for word usages, clustering these embeddings, and mapping between dictionary meaning entries and the resulting clusters. However, we make modifications to each component. The component-wise system comparison is presented in Figure 1.
  • For Subtask 2, unlike the baseline system, which requires costly model training for generating dictionary-like definitions for unmatched word usages, our system is training-free and does so by just prompting Large Language Models such as GPT-4 (Achiam et al., 2023) and LLaMA-3444https://llama.meta.com/llama3/. We provide the system comparison in Figure 2.
@@ 2 Related Work
focuses on the automatic identification of shifts in word meanings over time. For instance, the word ‘chill’ used to mean ‘cold’ for individuals growing up in the 60s, but for those in the 90s, it means ‘relaxed’. Many works proposed to detect meaning shifts by using static or contextualized embeddings (Eger and Mehler, 2016; Hamilton et al., 2016a, b; Martinc et al., 2020; Gonen et al., 2020; Kaiser et al., 2021; Montariol et al., 2021a; Teodorescu et al., 2022; Homskiy and Arefyev, 2022). Most work in LSCD has been done on an unsupervised task formulation (Schlechtweg et al., 2020) which neither involved a dictionary, nor providing interpretation or qualification of detected sense changes. While early work on static embeddings (Kim et al., 2014; Hamilton et al., 2016c) could qualify changes to a certain extent through nearest neighbors, it usually did not provide sense clusters in a dictionary-like manner. More recent work straightforwardly enables the induction of sense clusters through clustering of contextualized embeddings (Giulianelli et al., 2020; Kudisov and Arefyev, 2022; Montariol et al., 2021b; Arefyev and Bykov, 2021). More recently, Ma et al. (2024) presented a graph-based clustering approach to detect gained word senses with low frequency, and offered interpretability by visualizing cross-language semantic changes. The works by Giulianelli et al. (2023); Fedorova et al. (2024) offer new ways of interpretability such as automatically generating sense definitions for usages from clusters. For an overview of recent model architectures incl. clustering approaches, see Zamora-Reina et al. (2022).
@@ 4.1 Subtask 1
Lastly, to map between dictionary sense entries with unknown usages of a target word at a later time period, the baseline system proposes to align dictionary entries with the collective meaning of each cluster. In particular, for each cluster, the system chooses the embedding of the first-indexed usage of the target word in the AXOLOTL-24 datasets. as the collective meaning represented by that cluster. It then computes the cosine similarity between that word usage embedding and the embedding of each dictionary entry (i.e., sense gloss). If the similarity score surpasses a predefined threshold, then all the word usages within that cluster are said to be matching that dictionary entry.
@@
For each target word $w$, we denote $\mathcal{C}_{w}=\{c_{1},c_{2},\dots,c_{n}\}$ as a word cloud consisting of a set of $d$-dimensional embeddings. Each embedding represents a corpus usage of the target word, and $n$ denotes the number of word usages available in a given corpus that contain that target word. We aim to partition $\mathcal{C}_{w}$ into $m$ clusters. Each cluster contains a subset of $\mathcal{C}_{w}$ representing embeddings of word usages with similar meanings. Our clustering method is illustrated in Algorithm 1.
@@
Our clustering method is similar to the bottom-up agglomerative clustering Sibson (1973) but differs in that we use a neighbor-based metric [^7] For agglomerative clustering, the distance between two clusters is calculated as the average pairwise distance between usage pairs based on their embeddings. For us, each pairwise distance is calculated as the bipartite matching score over $k$-nearest neighbors of a word usage and those of another. to handle low-frequency clusters. The idea is the following: We start by treating each embedding as a separate cluster, and then iteratively merges two clusters when their centroids are of a distance smaller than the distance threshold $t_{sc}$ until no further pairs of such similar clusters can be found. Following Ma et al. (2024), we use a neighbor-based distance metric in the clustering process to compute distances between clusters. Both the distance threshold and the number of nearest neighbors are hyperparameters, which we tune on dev sets.
@@ 5 Experiments
For Subtask 1, the Adjusted Rand Index (ARI) (Hubert and Arabie, 1985) and macro-F1 score are the two evaluation metrics for reporting and comparing system performances. ARI calculates how much a pair of word usages from the predictions belong to the same sense ID (or different sense IDs) as they should, while the macro-F1 score computes the precision and recall of word usages for each sense ID and then averages these scores across all sense IDs. Note that novel senses are excluded when computing F1, meaning that mappings of word usages to novel senses unrecorded in dictionaries are not evaluated. On the contrary, mappings to both novel and old senses are evaluated when computing ARI.
@@ 6 Results
We made two submissions for Subtask 1, with minor difference between them. The only difference is that the second submission includes additional predictions for the unseen German language. Table 4 compares the results of our system and other teams. We see that our system, based on the unsupervised graph-based clustering approach, outperforms the unsupervised baseline system by a large margin in all the languages. We observe a big performance drop for the German language compared to other two languages. This might be due to the uncareful data preprocessing we did on the German test set: We find that quotes in German word usages are oftentimes escaped with double quotes, and we did not remove these additional quotes during the data loading process. This influences the outcome of word segmentation by m-BERT, potentially resulting in lower embedding quality. Lastly, our system performs poorly in terms of ARI on both Russian and German test sets, despite having better scores in macro-F1. The performance gap between F1 and ARI might attribute to the mismatch between the two metrics: old sense IDs are excluded when computing F1, whereas both old and novel sense IDs are included when computing ARI. This means unlike ARI, F1 would not penalize wrong mappings of word usages to novel sense IDs. Our further analysis confirms this hypothesis—we see that the performance gap becomes much smaller when ARI excludes novel sense IDs (cf. Table 6). We note that ARI is affected most in Russian, as the Russian test set contains many more novel senses (47%) compared to German (10.2%) and Finnish (5.8%), and the ARI result on those novel sense IDs is nearly zero.

### v2

@@ Abstract
There has been a surge of interest in computational modeling of semantic change. The foci of previous works are on detecting and interpreting word senses gained over time; however, it remains unclear whether the gained senses are covered by dictionaries. In this work, we aim to fill this research gap by comparing detected word senses with dictionary sense inventories in order to bridge between the communities of lexical semantic change detection and lexicography. We evaluate our system in the AXOLOTL-24 shared task for Finnish, Russian and German languages Fedorova et al. (2024b). Our system is fully unsupervised. It leverages a graph-based clustering approach to predict mappings between unknown word usages and dictionary entries for Subtask 1, and generates dictionary-like definitions for those novel word usages through the state-of-the-art Large Language Models such as GPT-4 and LLaMA-3 for Subtask 2. In Subtask 1, our system outperforms the baseline system by a large margin, and it offers interpretability for the mapping results by distinguishing between matched and unmatched (novel) word usages through our graph-based clustering approach. Our system ranks first in Finnish and German, and ranks second in Russian on the Subtask 2 test-phase leaderboard. These results show the potential of our system in managing dictionary entries, particularly for updating dictionaries to include novel sense entries. Our code and data are made publicly available [^1] https://github.com/xiaohemaikoo/axolotl24-ABDN-NLP.
@@ 1 Introduction
  • For Subtask 1, we keep the workflow of the AXOLOTL-24 baseline system unchanged, which includes three components: producing embeddings for word usages, clustering these embeddings, and mapping between dictionary meaning entries and the resulting clusters. However, we make modifications to each component. The component-wise system comparison is presented in Table 1.
  • For Subtask 2, unlike the baseline system, which requires costly model training for generating dictionary-like definitions for unmatched word usages, our system is training-free and does so by just prompting Large Language Models such as GPT-4 (Achiam et al., 2023) and LLaMA-3444https://llama.meta.com/llama3/. We provide the system comparison in Table 2.
@@ 2 Related Work
focuses on the automatic identification of shifts in word meanings over time. For instance, the word ‘chill’ used to mean ‘cold’ for individuals growing up in the 60s, but for those in the 90s, it means ‘relaxed’. Many works proposed to detect meaning shifts by using static or contextualized embeddings (Eger and Mehler, 2016; Hamilton et al., 2016a, b; Martinc et al., 2020; Gonen et al., 2020; Kaiser et al., 2021; Montariol et al., 2021a; Teodorescu et al., 2022; Homskiy and Arefyev, 2022). Most work in LSCD has been done on an unsupervised task formulation (Schlechtweg et al., 2020) which neither involved a dictionary, nor providing interpretation or qualification of detected sense changes. While early work on static embeddings (Kim et al., 2014; Hamilton et al., 2016c) could qualify changes to a certain extent through nearest neighbors, it usually did not provide sense clusters in a dictionary-like manner. More recent work straightforwardly enables the induction of sense clusters through clustering of contextualized embeddings (Giulianelli et al., 2020; Kudisov and Arefyev, 2022; Montariol et al., 2021b; Arefyev and Bykov, 2021). More recently, Ma et al. (2024) presented a graph-based clustering approach to detect gained word senses with low frequency, and offered interpretability by visualizing cross-language semantic changes. The works by Giulianelli et al. (2023); Fedorova et al. (2024a) offer new ways of interpretability such as automatically generating sense definitions for usages from clusters. For an overview of recent model architectures incl. clustering approaches, see Zamora-Reina et al. (2022).
@@ 4.1 Subtask 1
Lastly, to map between dictionary sense entries with unknown usages of a target word at a later time period, the baseline system proposes to align dictionary entries with the collective meaning of each cluster. In particular, for each cluster, the system chooses the embedding of the first-indexed usage of the target word in the AXOLOTL-24 datasets as the collective meaning represented by that cluster. It then computes the cosine similarity between that word usage embedding and the embedding of each dictionary entry (i.e., sense gloss). If the similarity score surpasses a predefined threshold, then all the word usages within that cluster are said to be matching that dictionary entry.
@@
For each target word $w$, we denote $\mathcal{C}_{w}=\{c_{1},c_{2},\dots,c_{n}\}$ as a word cloud consisting of a set of $d$-dimensional embeddings. Each embedding represents a corpus usage of the target word, and $n$ denotes the number of word usages available in a given corpus that contain that target word. We aim to partition $\mathcal{C}_{w}$ into $m$ clusters. Each cluster contains a subset of $\mathcal{C}_{w}$ representing embeddings of word usages with similar meanings. Our clustering method is illustrated in Algorithm 1. We choose our clustering over the baseline Affinity Propagation Frey and Dueck (2007) because target words in the AXOLOTL-24 datasets have 2-23 usages on average (c.f. Table 3), i.e., they only have low-frequency senses; in such a setup, our clustering largely outperforms Affinity Propagation (see Table 11 in Ma et al. (2024)). We present our clustering details in the following:
@@
Our clustering method is similar to the bottom-up agglomerative clustering Sibson (1973) but differs in that we use a neighbor-based metric [^7] For agglomerative clustering, the distance between two clusters is calculated as the average pairwise distance between usage pairs based on their embeddings. For us, each pairwise distance is calculated as the bipartite matching score over $k$-nearest neighbors of a word usage and those of another. to handle low-frequency clusters. The idea is the following: We start by treating each embedding as a separate cluster, and then iteratively merge two clusters when their centroids are of a distance smaller than the distance threshold $t_{sc}$ until no further pairs of such similar clusters can be found. Following Ma et al. (2024), we use a neighbor-based distance metric in the clustering process to compute distances between clusters. Both the distance threshold and the number of nearest neighbors are hyperparameters, which we tune on dev sets.
@@
Lastly, for mapping, we select the average usage embedding (i.e., cluster centroid) as the collective meaning of a cluster, and compare that embedding with dictionary entries. We choose the average embedding over the embedding of the first-indexed usage of a target word (see Table 1) because the first-indexed choice is almost random. We use the average embedding to eliminate such randomness.
@@ 5 Experiments
For Subtask 1, the Adjusted Rand Index (ARI) (Hubert and Arabie, 1985) and macro-F1 score are the two evaluation metrics for reporting and comparing system performances. ARI calculates how much a pair of word usages from the predictions belong to the same sense ID (or different sense IDs) as they should, while the macro-F1 score computes the precision and recall of word usages for each sense ID and then averages these scores across all sense IDs. Note that F1 only considers old senses in the “new” time period, meaning that mappings of word usages to novel senses are not evaluated. ARI considers both novel and old senses in the “new” time period.
@@ 6 Results
We made two submissions for Subtask 1, with minor difference between them. The only difference is that the second submission includes additional predictions for the unseen German language. Table 4 compares the results of our system and other teams. We see that our system, based on the unsupervised graph-based clustering approach, outperforms the unsupervised baseline system by a large margin in all the languages. We observe a big performance drop for the German language compared to other two languages. One of the reasons for this is due to historical data issues. Unlike the Russian and Finnish corpus usages—which have been carefully preprocessed by AXOLOTL-24 organizers, German usages are not cleaned up and contain spelling variations (e.g., nöthig instead of nötig), OCR errors, escaping double quotes and others. These issues would incur out-of-vocabulary tokens, potentially resulting in poor performance.
Lastly, our system performs poorly in terms of ARI on both Russian and German test sets, despite having better scores in macro-F1. The performance gap between F1 and ARI attributes to the scope mismatch between the two metrics: new sense IDs are excluded when computing F1, whereas both old and novel sense IDs are considered when computing ARI. This means unlike ARI, F1 would not penalize wrong prediction of novel sense IDs. As a result, although our system performs poorly for novel sense predictions in Russian (see the ARI_new result in Table 6), the F1 result (F1$=$0.570) is still quite high.
@@ 8 Conclusions
Acknowledgements
We thank the anonymous reviewers for their thoughtful feedback that greatly improved the texts. Dominik Schlechtweg has been funded by the research program ‘Change is Key!’ supported by Riksbankens Jubileumsfond (under reference number M21-0021).